OpenStudy (linn99123):

NEED HELP FAST WILL FAN AND GIVE MEEDAL Which set of integers is a Pythagorean triple? A. 3, 5, 6 B. 5, 12, 13 C. 15, 9, 17 D. 6, 8, 12

OpenStudy (anonymous):

I think it is B.

OpenStudy (anonymous):

B because it works out quite nicely. A2 + B2 = C2 (with C being the hypotenusethe longest side length) so 5^2 + 12^2 = 13^2 so 25 + 144 = 169 so 169 = 169 so it works :) cheers

OpenStudy (anonymous):

5²+12²=13² = 25+144=169 and so 169=169